OpenStudy (anonymous):

a computer sells for 995, which is marked up 35% of the selling price. the cost of the computer is?

OpenStudy (anonymous):

x + x*.35 = 995 1.35 x = 995 995/1.35 = x

OpenStudy (anonymous):

737.037

OpenStudy (anonymous):

answer = 995 + (995 * .35)